# Text-to-Model: Text-Conditioned Neural Network Diffusion for Train-Once-for-All Personalization

### v1

@@ 3.3 In-depth Analysis of Tina
Scaling studies for Tina. Scaling law was found for transformer-based foundation models that scaling the parameters, data, computes can bring intelligence emergence. In Figure 4 (a), we scale the parameters of Tina by changing the hidden sizes ranging from 32 (152M parameters) to 2048 (789M), and we test two sizes of p-Model. It is found that when Tina is small, it fails to generalize, especially when the p-Model has a higher parameter dimension. The intelligence emerges when scaling Tina at large sizes (e.g., 1024 or 2048 hidden sizes), but the scaling effect is saturated if reaching the upper bound performance of personalization. We also scale the input, also the generated, dimensions (i.e., p-Model sizes) and the training data in Figure 5. It is found that a larger input dimension is harder to learn and requires larger sizes of training data to converge and generalize. The generalization of Tina can benefit from larger training data, but it has diminishing marginal returns. Generally, larger p-Models, larger training samples, and larger model sizes make Tina reach higher p-Acc, and it demonstrates the increasing expressive power of Tina by scaling, which is consistent with previous DiT works [12, 16, 3]. The scaling property indicates the great potential of Tina for more complex and challenging text-to-model scenarios.
(a) Image as prompts.
(b) Different personalized classes.
(c) Descriptions as prompts.
Parameter inheritance. We verify whether Tina can benefit from pretrained parameters. We inherit the parameters from G.pt’s [16] checkpoints by the bert2BERT-like method [24]. From Figure 4 (b), it is found that parameter inheritance from pretrained models can help Tina to converge faster, but the final p-Accs are similar.
Training images as prompts. In the original design of Tina, the texts are used for the prompts encoded by the CLIP text encoder. We train a Tina with image prompts using CLIP image encoder, and the results are in Figure 4 (c). For each class, we randomly select one single image as the prompts. It is found that text-prompted Tina converges faster than the image-prompted, though the final p-Accs are similar. This is intuitive to understand since texts are known to have higher knowledge density than images [30, 17], that the class text has richer knowledge representations than a single image.
How Tina understands world knowledge I: natural language descriptions as prompts. In our implementation of Tina, we adopt a simple prompting that uses the class names as the text prompts. We verify whether Tina actually learns the knowledge in the case where the prompts are replaced by the natural language descriptions at test time. We generate the language descriptions of classes with the assistance of GPT-4 [31], and we make sure that the descriptions do not include the original class entities. The exemplars are in Table 4 of the appendix. From Figure 6 (c), the results reveal that Tina has zero-shot generalization ability when the prompts are unseen language descriptions, though the p-Accs are lower than the ones of the class-named prompts. It shows that Tina is not just memorizing the class names but also generalizing and understanding the knowledge behind the names and the nuances inherent in the text semantics.
@@
Settings 0% unseen tasks 20% unseen tasks 40% unseen tasks 60% unseen tasks 100% unseen tasks
TAPER-Mixer 60.27 51.94 42.48 31.45 0.0
@@
How Tina understands world knowledge II: generalization to unseen classes/entities. We divide the CIFAR-100 dataset into two disjoint shards of classes and train a Tina on one shard, then verify its generalization on the unseen classes of another shard. Results in Table 2 showcase that Tina has the intelligence to generalize on unseen classes, while TAPER-Mixer fails when meeting 100% unseen classes. As a generative model, Tina can understand the hyper-level world knowledge embedded in model parameters as well as text semantics and generate models for predicting unseen entities.
3.4 Ablation of Design Choices of Tina
We make an ablation study for different design choices of Tina. The ablated designs are the ones different from previous literature, such as our design of classifier augmentation, G.pt’s design of permutation augmentation [16], and TAPER’s design of merge text embedding as one [13]. The results are in Table 3. Our classifier augmentation can boost the performance even under small training datasets.
Table 3: Ablation study for different design choices of Tina.
Designs/Datasets Mini-Imagenet CIFAR-100 Caltech-101 Avg.
w/o classifier aug. 32.45 49.61 41.61 41.22
w/ permutation aug. 9.88 10.14 10.59 10.20
merge text embed. as one 10.04 10.35 10.78 10.39
Tina (completed) 53.31 67.14 59.27 59.91
4 Related Works
Diffusion models. Originating from non-equilibrium thermodynamics [32, 33], diffusion models have evolved significantly. DDPM and DDIM pioneered forward-and-reverse processes in text-to-image generation [9, 34]. Guided-based diffusion models [35] surpassed GAN-based methods in image generation quality. Subsequent models like GLIDE [36], Imagen [37], DALL·E 2 [2], and stable diffusion [4] further advanced image generation and art creation. The diffusion transformer (DiT) [12] introduced a scaling law, with OpenAI’s Sora [3] being a notable application in text-to-video generation, employing DiT architecture at a billion-scale.
Parameter generation. Learning to optimize explores neural networks learning update rules for others [38, 39, 40, 41]. Hypernetwork [42] is a meta learning approach that uses networks to modify neural network parameters, differing from our approach of mapping language space directly to parameter space. Hypernetworks are used in federated learning [43], few-shot learning [44], and model editing [45]. A concurrent work ModelGPT [46] customizes models by large language models and hypernetworks, while Tina uses conditional neural network diffusion for a different task—train-once-for-all personalization. Neural network diffusion [16, 15] is recently proposed to mimic optimization rules via diffusion for parameter generation, but previous works haven’t explored sufficient use cases of such techniques.
For more detailed related works (e.g., the works about personalization), please refer to Appendix B.
5 Discussions
Limitations. Despite the merits of Tina, it has some current limitations. One bottleneck is the input dimension; due to our computation limits, Tina currently supports lightweight models as inputs, and it requires huge computation resources to fully generate large models with millions of parameters. On the one hand, a larger input dimension needs exponentially larger Tina parameters, so more GPUs. On the other hand, a larger input dimension needs more data to converge or generalize, requiring more compute hours. As a remedy, we tried to train a variational autoencoder (VAE) for encoding the p-Model parameters into a low-dimension latent space as in [15], but the VAE cannot generalize, suggesting more advanced techniques are needed. Another limitation is the generality of Tina, that one single Tina cannot generate personalized models across different sizes and different modalities; in the future, large-scaling pretraining for Tina may be promising to reach this goal.
@@ A.1 Dataset Preparation
The Mini-ImageNet dataset [28] is a sub-dataset of ImageNet [27], which is widely used in few-shot learning. It selects 100 categories from ImageNet1K. The trainset contains 600 labeled images for each category, a total 60,000 images, and the testset contains 100 labeled images for each category, a total of 10,000 pieces.
@@
Each image in CIFAR-100 [14] has two labels: superclass and subclass. There are 500 training images and 100 testing images per subclass. CIFAR-100 has 20 superclasses, and each superclass has 5 subclasses.
@@
Caltech-101 [29] is an objects image dataset with 101 categories. Approximately 40 to 800 images per category, most categories have around 50 images, 8677 images in total. We divide it into a trainset and a testset according to the ratio of 8:2.
@@ A.2 Example of class description from GPT-4
For the word of each class, we use GPT-4 to provide a more detailed and standardized description and definition. Some examples are shown in Table 4.
@@ A.4 Detailed Implementations of Methods
TAPER-Mixer: We set up two base models and split the dataset into two shards based on the classification labels. Each base model is initialized using the parameters of the stage-one model and fine-tuned on one of the sharded datasets for 5 epochs. In stage 3, we use the class order of the p-Model in the trainset to train the mixer for 5 epochs, and during the testing phase, the mixer remains frozen.
@@ A.6 Environments and Resources
Appendix B Detailed Related Works
@@
The origin of diffusion models is the study of non-equilibrium thermodynamics [32, 33]. In recent years, DDPM [9] and DDIM [34] have refined diffusion models to a higher level by transforming the paradigm into forward-and-reverse processes in text-to-image generation. Later on, guided-based diffusion models [35] found a better architecture to improve the image generation quality that could beat the GAN-based methods [47, 48]. Then, GLIDE [36], Imagen [37], DALL·E 2 [2], and stable diffusion [4] emerged and flourished in the field of image generation and art creation. In the work of diffusion transformer (DiT) [12], the authors found that if the basic architecture of diffusion models is changed to transformers, the scaling law emerges, that scaling the number of parameters can reach the increasing quality of image generation. Based on DiT, in Feb 2024, OpenAI launched Sora [3], a text-to-video model that can understand and simulate the physical world in motion. In Sora, the DiT architecture is used and scaled to the billions level.
@@
The field of learning to optimize studies how one neural network can learn the update rules (gradients) for optimizing another network [38, 39, 40, 41]. Besides, the studies of hypernetworks [42] focus on how to directly output or modify neural networks’ parameters by a hypernetwork. Hypernetworks usually take models’ parameters as input and generate parameters [43, 45], which is different from our paper, which directly maps language space into the parameter space. Hypernetworks were used to generate local models for federated learning [43], edge-cloud collaboration, few-shot learning [44], and model editing [45]. A concurrent work ModelGPT [46] also uses text prompts to generate customized models by using large language models as task descriptors. However, ModelGPT didn’t target the train-once-for-all personalization scenario, and it uses conventional hypernetwork and meta learning methods while our Tina adopts novel conditional neural network diffusion. Recently, empowered by the strong expressiveness of diffusion models, neural network diffusion [16, 15] was proposed to mimic the optimization rule by diffusion for generating the model parameters. The first paper is G.pt [16], which uses DiT to learn to generate the model given a targeted loss or accuracy, and it mimics the optimization process while achieving faster inference compared with vanilla optimization. However, G.pt may have limited use cases; it can only generate the models for the training tasks (i.e., the in-distribution tasks in our paper’s terminology), and the accuracies are upper-bounded by the accuracies of checkpoint models in the training datasets. p-diff [15] formally formulates the neural network diffusion problem and proposes to diffuse and generate the batch normalization layers for better accuracies, but the improvement may be marginal, and the diffusion design is not conditioned. It also meets the dilemma of G.pt, which lacks a specific scenario and use case. Recently, GPD [18] uses the diffusion model for few-shot learning in smart city applications, which showcases the applications of neural network diffusion. However, GPD takes the smart city’s knowledge graphs as prompts and is tailored for the specific smart city application that cannot be easily extended to other fields. Our Tina takes language texts as prompts, which is more flexible and can be extended to a wider range of applications for the personalization of user demands.
@@
Instead of training a generic model to provide many users with the same model service, personalization of deep learning models acknowledges users’ characteristics and diversity and learns each a customized model. Personalization techniques were introduced in medical AI [49, 50, 51], recommendation systems [52, 53], large language models [54, 55], and especially federated learning [56, 57]. Personalized federated learning studies how to exploit the common knowledge of users and then use it to explore further personalization on users’ local datasets under privacy constraints [56], and techniques like proximal descent [58, 57], network decoupling [56, 59], and clustering [60, 61] are used. Recently, the scenario of train-once-for-all personalization [13] was proposed to bridge the gap between edge-side and server-side personalization. Train-once-for-all personalization aims to utilize server-side computation and generic models for fast and effective personalized adaptation to meet the edge users’ demands. The original method TAPER [13] finetunes the generic model into several base models and learns MLP-based hypernetworks as mixers to fuse the base models into the personalized one given users’ task descriptions. However, the MLP mixer has limited generalization capability, and it cannot be applied to unseen classes, whereas our Tina learns the text-to-model world knowledge and can be generalized to out-of-distribution samples, modalities, and domains.

### v2

@@ 3.3 In-depth Analysis of Tina
Scaling studies for Tina. Scaling law was found for transformer-based foundation models that scaling the parameters, data, computes can bring intelligence emergence. In Figure 4 (a), we scale the parameters of Tina by changing the hidden sizes ranging from 32 (152M parameters) to 2048 (789M), and we test two sizes of p-Model. It is found that when Tina is small, it fails to generalize, especially when the p-Model has a higher parameter dimension. The intelligence emerges when scaling Tina at large sizes (e.g., 1024 or 2048 hidden sizes), but the scaling effect is saturated if reaching the upper bound performance of personalization. We also scale the input, also the generated, dimensions (i.e., p-Model sizes) and the training data in Figure 5. It is found that a larger input dimension is harder to learn and requires larger sizes of training data to converge and generalize. The generalization of Tina can benefit from larger training data, but it has diminishing marginal returns. Generally, larger p-Models, larger training samples, and larger model sizes make Tina reach higher p-Acc, and it demonstrates the increasing expressive power of Tina by scaling, which is consistent with previous DiT works (Peebles & Xie, 2023; Peebles et al., 2022; OpenAI, 2024). The scaling property indicates the great potential of Tina for more complex and challenging text-to-model scenarios.
Parameter inheritance. We verify whether Tina can benefit from pretrained parameters. We inherit the parameters from G.pt’s (Peebles et al., 2022) checkpoints by the bert2BERT-like method (Chen et al., ). From Figure 4 (b), it is found that parameter inheritance from pretrained models can help Tina to converge faster, but the final p-Accs are similar.
Training images as prompts. In the original design of Tina, the texts are used for the prompts encoded by the CLIP text encoder. We train a Tina with image prompts using CLIP image encoder, and the results are in Figure 4 (c). For each class, we randomly select one single image as the prompts. It is found that text-prompted Tina converges faster than the image-prompted, though the final p-Accs are similar. This is intuitive to understand since texts are known to have higher knowledge density than images (Jia et al., 2021; Radford et al., 2021), that the class text has richer knowledge representations than a single image.
@@
Settings 0% unseen classes 20% unseen classes 40% unseen classes 60% unseen classes 100% unseen classes
TAPER 60.27 51.94 42.48 31.45 /
@@
How Tina understands world knowledge II: generalization to unseen classes/entities. We divide the CIFAR-100 dataset into two disjoint shards of classes and train a Tina on one shard, then verify its generalization on the unseen classes of another shard. Results in Table 3 showcase that Tina has the intelligence to generalize on unseen classes, while TAPER fails when meeting 100% unseen classes. As a generative model, Tina can understand the hyper-level world knowledge embedded in model parameters as well as text semantics and generate models for predicting unseen entities.
@@ A.1 Dataset Preparation
The Mini-ImageNet dataset (Vinyals et al., 2016) is a sub-dataset of ImageNet (Deng et al., 2009), which is widely used in few-shot learning. It selects 100 categories from ImageNet1K. The trainset contains 600 labeled images for each category, a total 60,000 images, and the testset contains 100 labeled images for each category, a total of 10,000 pieces.
@@
Each image in CIFAR-100 (Krizhevsky, 2009) has two labels: superclass and subclass. There are 500 training images and 100 testing images per subclass. CIFAR-100 has 20 superclasses, and each superclass has 5 subclasses.
@@
Caltech-101 (Fei-Fei et al., 2004) is an objects image dataset with 101 categories. Approximately 40 to 800 images per category, most categories have around 50 images, 8677 images in total. We divide it into a trainset and a testset according to the ratio of 8:2.
@@ A.2 Example of class description from GPT-4
For the word of each class, we use GPT-4 to provide a more detailed and standardized description and definition. Some examples are shown in Table 4. The prompts are:
"I will give you a list containing various nouns. Please add some short, accurate, and common descriptions to these nouns that can accurately define these nouns, and then return to me a JSON file where the key is the name and the value is the corresponding description. An example of the description is: "goblet": "a drinking glass with a base and stem", "anemones fish": "live associated with sea anemones", "chiffonier": "a tall elegant chest of drawers". The list to be processed is as follows:"
@@ A.4 Detailed Implementations of Methods
TAPER: We set up two base models and split the dataset into two shards based on the classification labels. Each base model is initialized using the parameters of the stage-one model and fine-tuned on one of the sharded datasets for 5 epochs. In stage 3, we use the class order of the p-Model in the trainset to train the mixer for 5 epochs, and during the testing phase, the mixer remains frozen.
@@ A.6 Environments and Resources
Appendix B More Results
Table 5: Analysis about whether Tina merely memorizes and reproduces parameters. The model is CNN, and the dataset is CIFAR-100. We verify Tina on OOD (unseen) tasks. Euclidean distances are calculated to reflect the parameter discrepancies directly. Also, we use model ensemble to verify whether the p-Models generated by Tina are functionally different and have diverse representations. Tina is conditioned on the class names as prompts during training. Here, we showcase two training tasks. “Finetune” refers to the oracle model finetuned on the target personalized dataset, “Tina ${}_{\text{name}}$” refers to Tina’s generated models during inference prompted on class names, “Tina ${}_{\text{des.}}$” refers to Tina’s generated models during inference prompted on class descriptions. Average accuracy (“Avg.”) refers to the average of individual accuracies. Ensemble accuracy (“Ensemble Acc.”) refers to ensembling the four models (1 “Finetune”, 2 “Tina ${}_{\text{name}}$”s, and 1 “Tina ${}_{\text{des.}}$”) during inference.
Individual Acc. Euclidean Distance
Finetune Tina name ${}_{\text{name 1}}$ Tina name ${}_{\text{name 2}}$ Tina ${}_{\text{des.}}$ Avg. Ensemble Acc. Tina ${}_{\text{name}}$-Finetune Tina ${}_{\text{des.}}$-Finetune Tina ${}_{\text{name}}$-Tina ${}_{\text{des.}}$
Task 1 75.3 74.9 74.9 58.9 71.0 76.2 4.11 11.41 10.72
Task 2 51.2 51.3 51.0 34.1 46.9 52.9 3.41 11.95 11.35
In Table 5, we additionally make an in-depth ablation study about whether Tina merely memorizes and reproduces parameters. The study includes the following aspects.
  • Euclidean Distances: It is found that the generated models have obvious Euclidean distances from each other and also from the fine-tuned models.
  • Ensemble Learning Ability: Ensemble learning often demonstrates higher accuracy than individual models, which can be indicative of the diversity in the internal representations of different neural networks, meaning that the manifold representations of the model parameters are not identical. Therefore, we make the generated models and the fine-tuned one ensemble to see whether it benefits. The results show that the ensemble accuracies are higher than the averaged accuracy and even higher than the best individual accuracy.
  • Taking the above experimental results into consideration, it is evident that Tina is not merely memorizing parameters but generalizing.
B.1 Ablation of Design Choices of Tina
Table 6: Ablation study for different design choices of Tina.
Designs/Datasets Mini-Imagenet CIFAR-100 Caltech-101 Avg.
w/o classifier aug. 32.45 49.61 41.61 41.22
w/ permutation aug. 9.88 10.14 10.59 10.20
merge text embed. as one 10.04 10.35 10.78 10.39
Tina (completed) 53.31 67.14 59.27 59.91
We make an ablation study for different design choices of Tina. The ablated designs are the ones different from previous literature, such as our design of classifier augmentation, G.pt’s design of permutation augmentation (Peebles et al., 2022), and TAPER’s design of merge text embedding as one (Chen et al., 2023). The results are in Table 6. Our classifier augmentation can boost the performance even under small training datasets.
Table 7: Ablation study of Tina on the impact of text prompts. The model is CNN and the dataset is CIFAR-100.
Testing Prompt
Training Prompt Class name Description
ID OOD ID OOD
Class name 67.27 67.21 46.93 46.77
Description 42.79 42.58 67.29 67.08
Ablation of text prompts. We have made an in-depth ablation study on the impact of text prompts, as in Table 7. It is found that if training and testing use the same kind of text prompts, the performances are similar regardless of class-name prompting or description prompting. However, if the prompt strategies are different in training and testing, the results will degrade, and training in class name prompts has better transferability and generalization.
Analysis about whether Tina merely memorizes and reproduces parameters. In Table 5 of Appendix, we additionally make an in-depth ablation study about whether Tina merely memorizes and reproduces parameters. We use Euclidean distances of parameters and ensemble learning ability to verify, and the results show that Tina is not merely memorizing but generalizing.
How Tina understands world knowledge I: natural language descriptions as prompts. In our implementation of Tina, we adopt a simple prompting that uses the class names as the text prompts. We verify whether Tina actually learns the knowledge in the case where the prompts are replaced by the natural language descriptions at test time. We generate the language descriptions of classes with the assistance of GPT-4 (OpenAI & the co authors, 2024), and we make sure that the descriptions do not include the original class entities. The exemplars are in Table 4 of the appendix. From Figure 6 (c), the results reveal that Tina has zero-shot generalization ability when the prompts are unseen language descriptions, though the p-Accs are lower than the ones of the class-named prompts. It shows that Tina is not just memorizing the class names but also generalizing and understanding the knowledge behind the names and the nuances inherent in the text semantics.
(a) Image as prompts.
(b) Different personalized classes.
(c) Descriptions as prompts.
Appendix C Related Works
Diffusion models. Originating from non-equilibrium thermodynamics (Jarzynski, 1997; Sohl-Dickstein et al., 2015), diffusion models have evolved significantly. DDPM and DDIM pioneered forward-and-reverse processes in text-to-image generation (Nichol & Dhariwal, 2021; Song et al., 2020). Guided-based diffusion models (Dhariwal & Nichol, 2021) surpassed GAN-based methods in image generation quality. Subsequent models like GLIDE (Nichol et al., 2021), Imagen (Saharia et al., 2022), DALL·E 2 (Ramesh et al., 2022), and stable diffusion (Rombach et al., 2022) further advanced image generation and art creation. The diffusion transformer (DiT) (Peebles & Xie, 2023) introduced a scaling law, with OpenAI’s Sora (OpenAI, 2024) being a notable application in text-to-video generation, employing DiT architecture at a billion-scale.
Parameter generation. Learning to optimize explores neural networks learning update rules for others (Andrychowicz et al., 2016; Amos, 2022; Metz et al., 2022; Chandra et al., 2022). Hypernetwork (Ha et al., 2016) is a meta learning approach that uses networks to modify neural network parameters, differing from our approach of mapping language space directly to parameter space. Hypernetworks are used in federated learning (Shamsian et al., 2021), few-shot learning (Zhmoginov et al., 2022), and model editing (Mitchell et al., 2022). A concurrent work ModelGPT (Tang et al., 2024) customizes models by large language models and hypernetworks, while Tina uses conditional neural network diffusion for a different task—train-once-for-all personalization. Neural network diffusion (Peebles et al., 2022; Wang et al., 2024) is recently proposed to mimic optimization rules via diffusion for parameter generation, but previous works haven’t explored sufficient use cases of such techniques.
@@ Appendix C Related Works
For more detailed related works (e.g., the works about personalization), please refer to Appendix E.
Appendix D Discussions
Limitations. Despite the merits of Tina, it has some current limitations. One bottleneck is the input dimension; due to our computation limits, Tina currently supports lightweight models as inputs, and it requires huge computation resources to fully generate large models with millions of parameters. On the one hand, a larger input dimension needs exponentially larger Tina parameters, so more GPUs. On the other hand, a larger input dimension needs more data to converge or generalize, requiring more compute hours. As a remedy, we tried to train a variational autoencoder (VAE) for encoding the p-Model parameters into a low-dimension latent space as in Wang et al. (2024), but the VAE cannot generalize, suggesting more advanced techniques are needed. Another limitation is the generality of Tina, that one single Tina cannot generate personalized models across different sizes and different modalities; in the future, large-scaling pretraining for Tina may be promising to reach this goal.
Appendix E Detailed Related Works
@@
The origin of diffusion models is the study of non-equilibrium thermodynamics (Jarzynski, 1997; Sohl-Dickstein et al., 2015). In recent years, DDPM (Nichol & Dhariwal, 2021) and DDIM (Song et al., 2020) have refined diffusion models to a higher level by transforming the paradigm into forward-and-reverse processes in text-to-image generation. Later on, guided-based diffusion models (Dhariwal & Nichol, 2021) found a better architecture to improve the image generation quality that could beat the GAN-based methods (Goodfellow et al., 2014; 2020). Then, GLIDE (Nichol et al., 2021), Imagen (Saharia et al., 2022), DALL·E 2 (Ramesh et al., 2022), and stable diffusion (Rombach et al., 2022) emerged and flourished in the field of image generation and art creation. In the work of diffusion transformer (DiT) (Peebles & Xie, 2023), the authors found that if the basic architecture of diffusion models is changed to transformers, the scaling law emerges, that scaling the number of parameters can reach the increasing quality of image generation. Based on DiT, in Feb 2024, OpenAI launched Sora (OpenAI, 2024), a text-to-video model that can understand and simulate the physical world in motion. In Sora, the DiT architecture is used and scaled to the billions level.
@@
The field of learning to optimize studies how one neural network can learn the update rules (gradients) for optimizing another network (Andrychowicz et al., 2016; Amos, 2022; Metz et al., 2022; Chandra et al., 2022). Besides, the studies of hypernetworks (Ha et al., 2016) focus on how to directly output or modify neural networks’ parameters by a hypernetwork. Hypernetworks usually take models’ parameters as input and generate parameters (Shamsian et al., 2021; Mitchell et al., 2022), which is different from our paper, which directly maps language space into the parameter space. Hypernetworks were used to generate local models for federated learning (Shamsian et al., 2021), edge-cloud collaboration, few-shot learning (Zhmoginov et al., 2022), and model editing (Mitchell et al., 2022). A concurrent work ModelGPT (Tang et al., 2024) also uses text prompts to generate customized models by using large language models as task descriptors. However, ModelGPT didn’t target the train-once-for-all personalization scenario, and it uses conventional hypernetwork and meta learning methods while our Tina adopts novel conditional neural network diffusion. Recently, empowered by the strong expressiveness of diffusion models, neural network diffusion (Peebles et al., 2022; Wang et al., 2024) was proposed to mimic the optimization rule by diffusion for generating the model parameters. The first paper is G.pt (Peebles et al., 2022), which uses DiT to learn to generate the model given a targeted loss or accuracy, and it mimics the optimization process while achieving faster inference compared with vanilla optimization. However, G.pt may have limited use cases; it can only generate the models for the training tasks (i.e., the in-distribution tasks in our paper’s terminology), and the accuracies are upper-bounded by the accuracies of checkpoint models in the training datasets. p-diff (Wang et al., 2024) formally formulates the neural network diffusion problem and proposes to diffuse and generate the batch normalization layers for better accuracies, but the improvement may be marginal, and the diffusion design is not conditioned. It also meets the dilemma of G.pt, which lacks a specific scenario and use case. Recently, GPD (Yuan et al., 2024) uses the diffusion model for few-shot learning in smart city applications, which showcases the applications of neural network diffusion. However, GPD takes the smart city’s knowledge graphs as prompts and is tailored for the specific smart city application that cannot be easily extended to other fields. Our Tina takes language texts as prompts, which is more flexible and can be extended to a wider range of applications for the personalization of user demands.
@@
Instead of training a generic model to provide many users with the same model service, personalization of deep learning models acknowledges users’ characteristics and diversity and learns each a customized model. Personalization techniques were introduced in medical AI (Goecks et al., 2020; Awwalu et al., 2015; Li et al., 2022b), recommendation systems (Choi et al., 2006; Cui et al., 2020), large language models (Kirk et al., 2024; Li et al., 2024a), and especially federated learning (Chen & Chao, 2022; Li et al., 2021). Personalized federated learning studies how to exploit the common knowledge of users and then use it to explore further personalization on users’ local datasets under privacy constraints (Chen & Chao, 2022), and techniques like proximal descent (Li et al., 2020; 2021), network decoupling (Chen & Chao, 2022; Gao et al., 2023), and clustering (Ghosh et al., 2020; Li et al., 2022a) are used. Recently, the scenario of train-once-for-all personalization (Chen et al., 2023) was proposed to bridge the gap between edge-side and server-side personalization. Train-once-for-all personalization aims to utilize server-side computation and generic models for fast and effective personalized adaptation to meet the edge users’ demands. The original method TAPER (Chen et al., 2023) finetunes the generic model into several base models and learns MLP-based hypernetworks as mixers to fuse the base models into the personalized one given users’ task descriptions. However, the MLP mixer has limited generalization capability, and it cannot be applied to unseen classes, whereas our Tina learns the text-to-model world knowledge and can be generalized to out-of-distribution samples, modalities, and domains.